1) What is the expected value of a single roll of the die? 2) What is...

1)

What is the expected value of a single roll of the die?

2)

What is the probability of rolling a 3.5?

3)

A study of a new cholesterol drug measures the decrease in LDL (the "bad" cholesterol) of a group of 5 of patients. The recorded changes in cholesterol were -20.4, -18.0, -7, -2.8, -3.2. What is the expected value of the drop in cholesterol for this drug?

Question 3 options:

The mean of the five values shown above.

Take the probability of each of the five values and multiply by the cholesterol value. Then add them.

Unable to determine. We would need to examine a density curve.

4)

A study of a new cholesterol drug measures the decrease in LDL (the "bad" cholesterol) of a group of 500 of patients. The drop in cholesterol followed an approximately Normal distribution: N(-20.0, 2). What is the probability that a person chosen at random will have a cholesterol drop of between -22 and -18?

Question 4 options:

About 16%

About 32%

About 68%

Unable to determine.
Answer #1

x = 1, 2, 3, 4, 5, 6 . P[X] is as follows in table.. - E (X) = Pcm) x + 4x 1 I + 2x1 + 3x 6 + 5x + 6xl f = 1 + 2 + 3 + 4 + sprob. of rollo a 3.5 It is not possible to rollang 3-s. Elx ) = -2 0 .4 + 1-7)+(-2.8)+ (-3.2) +6-18.0 3 -5144 S - 10.28 | Ans